Find the value(s) of x for which f(x)=g(x)
f(x)=x"squared"+10x-14, g(x)=9x-8
so far i have:
x"squared"+10x-14=9x-8

Answer by Nate(3500)

x^2 + 10x - 14 = 9x - 8
x^2 + x = 6
(x + 0.5)^2 = 6.25
x + 0.5 = +- 2.5
x = -0.5 +- 2.5
x = -3 and x = 2